JosephGunnels:

A circle that has (3,-1) and (7,-5) as endpoints of a diameter. Use the given information to write the standard equation of the circle.

JosephGunnels:

plz i need someones help

Vocaloid:

you need two pieces of information to write the equation for a circle - the radius r - the center (h,k) standard equation is (x-h)^2 + (y-k)^2 = r^2 you're given the endpoints of the diameter. to find the radius, use the distance formula on the two points to find the diameter, and then divide that by 2 to get the radius. that'll be your r value to find the center, use the midpoint formula on the two points to get the midpoint (h,k) plug h, k, and r into the standard equation (x-h)^2 + (y-k)^2 = r^2

JosephGunnels:

im rlly confused how to put it down cs my geometry teacher gets the questions wrong himself so i dont rlly know to put it down as

JosephGunnels:

can you help me?

Vocaloid:

so you know the general standard equation for a circle now, (x-h)^2 + (y-k)^2 = r^2, and you need to find the h, k, and r values based on the information I've given you. are you getting stuck on some specific part?

JosephGunnels:

all of it cs my teacher hasnt even taught us this

Vocaloid:

to describe a circle, the standard equation is (x-h)^2 + (y-k)^2 = r^2, where r is the radius, and (h,k) is the center. h is the x-value of the center, k is the y-value of the center. the problem is asking you to write the equation of a circle. its diameter has endpoints (3,-1) and (7,-5). to find the length of the diameter, we need the distance between those two points. to do that, we use the distance formula \[d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\] plug in your two points (3,-1) and (7,-5) into this formula to find the distance. pick one of the points to be (x1,y1) and the other one to be (x2,y2)

JosephGunnels:

ok i did the distance formula and i got a decimal is tht correct?

Vocaloid:

what number did you get?

JosephGunnels:

i got the distance formula as 6.32455532

Vocaloid:

I think you made some arithmetic error somewhere \[d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\] plugging in (3,-1) and (7,-5) \[d=\sqrt{(7-3)^{2}+(-5-(-1))^{2}}\]=\[\sqrt{4^2+(-4)^{2}}=\sqrt{32}=\sqrt{4*4*2}=4\sqrt{2}\] let's leave it as a radical for now. if the diameter is 4sqrt(2), that means our radius is one half of that, our 2sqrt(2) so that's our r value, let's move onto the center, which is thankfully a bit easier

JosephGunnels:

how do you get \[4\sqrt{2}\]

Vocaloid:

\[d=\sqrt{(7-3)^{2}+(-5-(-1))^{2}}\]=\[\sqrt{4^2+(-4)^{2}}=\sqrt{32}=\sqrt{4*4*2}=4\sqrt{2}\] 32 is equal to 4 * 4 * 2 the square root of (4 * 4) is 4 so sqrt(4 * 4 * 2) becomes 4*sqrt(2)

JosephGunnels:

ok

JosephGunnels:

so whts do you do next?

Vocaloid:

now that we have the radius r, we can find the center of the circle the center of a circle is the midpoint of the diameter our diameter has endpoints (3,-1) and (7,-5), so we need to find the midpoint of these two points midpoint formula: \[(\frac{ x_{1}+x_{2} }{ 2 },\frac{ y_{1}+y_{2} }{ 2 })\] basically, add the x-coordinates of (3,-1) and (7,-5), divide by 2, that's the x-coordinate of the midpoint repeat with the y-coordinates to get the y-coordinate of the midpoint lmk what you get or if you're still stuck

JosephGunnels:

ok

JosephGunnels:

question do you subtract y1 and y2 together or do you subtract since they are bothe a negative?

Vocaloid:

you add the two values, regardless of their sign -1 + (-5) is the same as -1 - 5 = -6

JosephGunnels:

ok so it would be -6 divided by 2?

Vocaloid:

yes

JosephGunnels:

ok and after you got them added you divide them?

Vocaloid:

yes, so -6/2 = -3 as your y-coordinate

JosephGunnels:

and 5 as my x- coordinate

JosephGunnels:

so wht do i do next?

Vocaloid:

good, so (5,-3) is your midpoint and thus the center of your circle putting everything together: r = 2sqrt(2), (h,k) = (5,-3) standard equation is (x-h)^2 + (y-k)^2 = r^2 so plugging in the r, h, and k values \[(x-5)^{2}+(y-(-3))^{2}=(2\sqrt{2})^2\] which simplifies to \[(x-5)^{2}+(y+3)^{2}=8\]as our standard form equation

JosephGunnels:

THANK YOU SO MUCH!